Suppose that you hold a piece of land in the City of London that you may want to sell in one year. As a U.S. resident, you are concerned with the dollar value of the land. Assume that, if the British economy booms in the future, the land will be worth £1,898 and one British pound will be worth $1.43. If the British economy slows down, on the other hand, the land will be worth less, i.e., £1,351, but the pound will be stronger, i.e., $1.62/£. You feel that the British economy will experience a boom with a 60% probability and a slow-down with a 40% probability.

(b) Compute the variance of the dollar value of your property that is attributable to the exchange rate uncertainty.
Answer #1

gled Answer: Let us compute the neces cessary parametervalues ; ECP)=(66X$2,7)+(64)42,189 = $1628.4+876 =$2504.4 =$2504. FCs)وع b=COWAYway (8 =-23 -23.94/ 0.008 --£2,758 WC get a negative Exposure fis the pound gety strogor (weaker) the dollar, the d